OpenStudy (anonymous):

4y-12=-4 solve for y

OpenStudy (anonymous):

im not good at this kind of math :\

OpenStudy (anonymous):

4y-12=-4 +12 +12 4y=8 Divide by 4 y=2

OpenStudy (shaik0124):

4y=-4+12 4y=-8 y=-2

OpenStudy (shaik0124):

sorry it's+2

OpenStudy (shaik0124):

4y=8 y=2

OpenStudy (anonymous):

thank you all like i said im not good at this kind of math

OpenStudy (e.cociuba):

Simplifying 4y + -12 = -4 Reorder the terms: -12 + 4y = -4 Solving -12 + 4y = -4 Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '12' to each side of the equation. -12 + 12 + 4y = -4 + 12 Combine like terms: -12 + 12 = 0 0 + 4y = -4 + 12 4y = -4 + 12 Combine like terms: -4 + 12 = 8 4y = 8 Divide each side by '4'. y = 2 Simplifying y = 2 SORCE: http://www.geteasysolution.com/4y-12=-4